# Adaptive Learning Guided by Bias-Noise-Alignment Diagnostics

### v1

@@ Abstract
Learning systems deployed in nonstationary and safety-critical environments often suffer from instability, slow convergence, or brittle adaptation when learning dynamics evolve over time. While modern optimization, reinforcement learning, and meta-learning methods adapt to gradient statistics, they largely ignore the temporal structure of the error signal itself. This paper proposes a diagnostic-driven adaptive learning framework that explicitly models error evolution through a principled decomposition into bias, capturing persistent drift; noise, capturing stochastic variability; and alignment, capturing repeated directional excitation leading to overshoot. These diagnostics are computed online from lightweight statistics of loss or temporal-difference error trajectories and are independent of model architecture or task domain. We show that the proposed bias–noise–alignment decomposition provides a unifying control backbone for supervised optimization, actor–critic reinforcement learning, and learned optimizers. Building on this framework, we derive diagnostic-driven instantiations including a stabilized supervised optimizer, a diagnostic-regulated actor–critic scheme, and a diagnostic-conditioned learned optimizer. Under standard smoothness assumptions, we establish bounded effective updates and stability properties for all cases. Representative diagnostic illustrations in actor–critic learning highlight how the proposed signals modulate adaptation in response to temporal-difference error structure. Overall, this work elevates error evolution to a first-class object in adaptive learning and provides an interpretable, lightweight foundation for reliable learning in dynamic environments.
@@
adaptive learning; optimization; reinforcement learning; learned optimizers; nonstationarity; stability; interpretability.
@@ 1 Introduction
Modern learning systems are increasingly deployed in environments that are nonstationary, noisy, and safety-critical, where the statistical properties of data and objectives evolve over time. In such settings, learning dynamics are often dominated not only by stochastic gradients but also by structured changes in operating conditions, feedback reliability, and temporal dependencies. Examples include sequential prediction under regime shifts, control under uncertain dynamics, and continual or task-adaptive learning. Despite remarkable progress in optimization and reinforcement learning, ensuring stable, interpretable, and reliable adaptation under these conditions remains a fundamental challenge.
@@
Most widely used optimization algorithms adapt learning rates based on instantaneous or exponentially averaged gradient statistics. Methods such as Adam and AdamW [kingma2014adam, loshchilov2017decoupled], large-batch optimizers such as LAMB [you2019large], sign-based methods such as Lion [chen2023lion], and second-order approximations such as Sophia [liu2023sophia] have proven effective across a wide range of applications. However, these methods primarily react to gradient magnitude and variance, without explicitly accounting for how the error signal itself evolves during training. In nonstationary settings, this limitation can lead to unstable convergence, oscillatory behavior, or overly conservative adaptation, particularly in long-horizon sequence models and control-oriented learning problems. Related challenges arise in reinforcement learning, where training stability is strongly influenced by the structure of temporal-difference (TD) errors. Stabilization techniques such as trust-region methods [schulman2015trust], proximal policy optimization (PPO) [schulman2017proximal], entropy regularization, and advantage estimation reduce variance and catastrophic updates, but they do not explicitly distinguish between systematic error drift and stochastic fluctuations in the learning signal. As a result, policy learning remains highly sensitive to hyperparameters, noise, and implementation details, particularly under changing dynamics or partial observability [henderson2018deep].
@@
Meta-learning and learned optimizers provide another avenue for adaptation by learning update rules or initializations that generalize across tasks. Approaches such as MAML [finn2017model], Meta-SGD [li2017meta], Reptile, and recurrent learned optimizers [andrychowicz2016learning] demonstrate that optimization itself can be learned from data. While powerful, these methods typically condition on gradients and internal hidden states, offering limited interpretability of why or when adaptation occurs. Moreover, without explicit mechanisms to assess the reliability of the error signal, learned optimizers may still exhibit brittle behavior under unseen or highly nonstationary conditions. A common thread across supervised learning, reinforcement learning, and meta-learning is that adaptation decisions are largely driven by gradient-based signals, while the temporal structure of the error signal is treated as incidental. In contrast, biological and human motor learning exhibit a qualitatively different behavior, adaptation is modulated based on perceived error trends, variability, and repeated overshoot rather than instantaneous feedback alone. Persistent drift triggers corrective action, high variability leads to cautious updates, and repeated overshoot prompts attenuation of movement amplitude. These observations suggest that error evolution contains interpretable structure that can serve as a principled control signal for adaptive learning.
@@
Motivated by this perspective, we propose a diagnostic-driven adaptive learning framework that explicitly models error dynamics through a simple yet expressive decomposition into three components: bias, capturing persistent error drift; noise, capturing stochastic variability around the trend; and alignment, capturing systematic overshoot arising from repeated update directions. These diagnostics are computed online from lightweight statistics of loss or TD-error evolution and are independent of model architecture or task domain. Rather than replacing gradient-based learning, the proposed diagnostics complement gradients by regulating update magnitude, direction, and exploration based on the reliability and structure of the learning signal. The central contribution of this work is to demonstrate that a bias–noise–alignment decomposition of error dynamics provides a unifying backbone for adaptive learning across multiple paradigms. We show that the same diagnostic signals can be systematically reused to stabilize supervised optimization, regulate actor–critic reinforcement learning, and condition learned optimizers for rapid task adaptation. By explicitly elevating error evolution to a first-class object, this framework bridges ideas from optimization, control, and meta-learning within a single interpretable perspective. Under standard smoothness assumptions, we establish bounded effective updates and descent-style stability guarantees for diagnostic-driven adaptation. The proposed approach does not replace existing optimizers or control algorithms; rather, it enhances their robustness when learning dynamics become biased, noisy, or nonstationary. The major contributions of this work are summarized below.
@@
  • We introduce a lightweight bias–noise–alignment decomposition of error dynamics that is model-agnostic and computable online from loss or TD-error trajectories.
@@
  • We instantiate a single diagnostic framework across supervised learning, actor–critic reinforcement learning, and meta-learned optimizers, demonstrating its role as a unifying mechanism for adaptive learning.
@@ 2 Review of Literature
This work connects adaptive optimization, reinforcement learning stabilization, and meta-learning. We briefly review relevant literature in these areas and highlight the limitations that motivate a diagnostic-driven perspective on adaptive learning.
@@ 2.1 Adaptive Optimization Methods
Stochastic gradient-based optimization remains the backbone of modern machine learning. Classical methods such as stochastic gradient descent (SGD) with momentum adapt update directions using first-order information, while adaptive methods such as Adam and AdamW [kingma2014adam, loshchilov2017decoupled] further normalize updates using gradient moments. These approaches provide robustness to scale differences across parameters and have become standard in training deep networks. Extensions such as AdaFactor [shazeer2018adafactor], LAMB [you2019large], and Lion [chen2023lion] target improved scalability, memory efficiency, or stability in large-batch and large-model regimes. More recently, sharpness-aware optimization methods such as SAM [foret2020sharpness] and its variants have been proposed to improve generalization by explicitly seeking flat minima. While these methods introduce an implicit sensitivity to local loss geometry, they typically rely on inner maximization steps or gradient perturbations and do not explicitly model the temporal evolution of the error signal. Second-order or quasi-second-order approaches, including Sophia [liu2023sophia], approximate curvature information to improve convergence speed, but remain primarily reactive to instantaneous gradient statistics. A common limitation across these methods is that adaptation decisions are driven by gradient magnitude, variance, or curvature proxies, rather than by structured information about how the error evolves over time. As a result, these optimizers do not explicitly distinguish between persistent error drift, stochastic fluctuations, or repeated overshoot, which can lead to instability or overly conservative updates under nonstationary learning conditions.
@@ 2.2 Stability in Reinforcement Learning
Reinforcement learning (RL) introduces additional challenges due to delayed rewards, bootstrapping, and nonstationary data distributions induced by policy updates  [dulac2021challenges]. Actor–critic methods mitigate variance by learning a value function, while advantage estimation techniques further stabilize gradient estimates. Trust-region approaches such as TRPO [schulman2015trust] constrain policy updates to avoid destructive changes, and PPO [schulman2017proximal] provides a computationally efficient approximation through clipped objectives. Entropy regularization is commonly used to encourage exploration and prevent premature convergence. However, entropy coefficients are typically fixed or scheduled heuristically, without explicit consideration of the structure of the TD error. Empirical studies have shown that deep RL methods remain highly sensitive to hyperparameters, random seeds, and implementation details [henderson2018deep], particularly under noisy or changing dynamics. While existing stabilization techniques limit update magnitude or variance, they do not explicitly distinguish whether instability arises from systematic bias in TD errors or from stochastic noise. Consequently, learning algorithms may respond uniformly to fundamentally different error regimes, motivating the need for diagnostic signals that can disentangle these effects.
@@ 2.3 Meta-Learning and Learned Optimizers
Meta-learning aims to enable rapid adaptation across tasks by learning how to learn. Model-agnostic meta-learning (MAML) [finn2017model] focuses on learning an initialization that supports fast task-specific adaptation using standard gradient descent. Extensions such as Meta-SGD [li2017meta] jointly learn initialization and per-parameter learning rates, improving flexibility and adaptation speed. An alternative line of work treats the optimization process itself as a learning problem [chen2022learning]. Learned optimizers parameterize update rules using neural networks and are trained through meta-gradient descent across tasks [andrychowicz2016learning]. Recurrent and coordinate-wise learned optimizers have demonstrated impressive data efficiency in few-shot and continual learning settings. Despite these advances, learned optimizers often rely on opaque internal states and gradient inputs, making it difficult to interpret or reason about their adaptation behavior. Moreover, without explicit mechanisms to assess the reliability of the learning signal, learned optimizers may generalize poorly to tasks with substantially different noise characteristics or nonstationary dynamics. This limitation suggests that conditioning learned updates on interpretable diagnostics of error structure may improve robustness and transparency.
@@ 2.4 Human-Inspired and Control-Theoretic Perspectives
A broad class of learning algorithms has been motivated by principles drawn from human cognition, motor learning, and classical control theory. Representative examples include gain scheduling and adaptive step-size control in optimization, uncertainty-aware exploration and entropy modulation in reinforcement learning, and curriculum or experience-driven adaptation in meta-learning. These approaches have demonstrated empirical benefits in stabilizing training and improving convergence in complex, nonstationary environments. Despite their success, most existing human-inspired methods incorporate adaptivity implicitly or heuristically. In practice, adaptation is often driven by gradient magnitude, loss curvature, or hand-designed schedules, without an explicit characterization of the reason behind the unstable learning. From a control-theoretic perspective, however, instability is rarely monolithic. Persistent bias such as systematic drift, stochastic noise such as measurement or process variability, and repeated directional excitation such as overshoot or resonance are treated as distinct phenomena, each requiring different compensatory mechanisms. When these effects are conflated, corrective actions may be either overly aggressive or excessively conservative. A similar limitation is observed across adaptive optimization, reinforcement learning stabilization, and meta-learning. Although these methods provide powerful tools for learning in high-dimensional and uncertain settings, they remain largely gradient-centric and do not explicitly model the temporal structure of error evolution. In particular, they do not distinguish between systematic and stochastic sources of instability, nor do they provide an interpretable signal for modulating update magnitude, direction, or exploration based on error reliability.
The present work addresses this gap by introducing a diagnostic decomposition of error dynamics into bias, noise, and alignment components. Inspired by both human motor learning and control-theoretic reasoning, these diagnostics explicitly separate persistent drift, stochastic variability, and directional overshoot phenomena that humans naturally compensate for by adjusting step size, movement direction, and caution under uncertainty. Rather than serving as ad hoc heuristics, the proposed diagnostics form a unified and interpretable control signal that regulates learning across supervised optimization, actor-critic reinforcement learning, and meta-learned optimizers. In this sense, the framework operationalizes human-inspired adaptation through a principled, task-agnostic representation of error evolution.
@@ 3 Diagnostic Decomposition of Error Dynamics
This section discusses the diagnostic framework used in this study to explicitly model the temporal evolution of error during learning. Rather than relying solely on instantaneous gradients, we extract interpretable signals from the dynamics of the loss or TD error and use them as lightweight controls for adaptive learning. The proposed decomposition separates error dynamics into three complementary components namely, bias, noise, and alignment. These diagnostics are general, model-agnostic, and applicable across supervised learning, reinforcement learning, and meta-learning. The details of each mechanism are discussed in the section below.
@@ 3.1 Error Evolution Signal
Let $\ell_{t}$ denote a scalar learning signal evaluated at iteration $t$. In supervised learning, $\ell_{t}$ may correspond to a training or validation loss, while in reinforcement learning it may represent a TD error or a batch-level surrogate thereof. Rather than treating $\ell_{t}$ as an isolated quantity, we focus on its temporal evolution. We define the incremental error signal which captures whether learning progress is improving, stagnating, or degrading over time. Unlike raw gradients, $\Delta\ell_{t}$ directly reflects changes in performance and aggregates the effects of optimization, stochasticity, and model mismatch.

### v2

@@ Abstract
Learning systems deployed in nonstationary and safety-critical environments often suffer from instability, slow convergence, or brittle adaptation when learning dynamics evolve over time. While modern optimization, reinforcement learning, and meta-learning methods adapt to gradient statistics, they largely ignore the temporal structure of the error signal itself. This paper proposes a diagnostic-driven adaptive learning framework that explicitly models error evolution through a principled decomposition into bias, capturing persistent drift; noise, capturing stochastic variability; and alignment, capturing repeated directional excitation leading to overshoot. These diagnostics are computed online from lightweight statistics of loss or temporal-difference (TD) error trajectories and are independent of model architecture or task domain. We show that the proposed bias-noise-alignment decomposition provides a unifying control backbone for supervised optimization, actor-critic reinforcement learning, and learned optimizers. Within this framework, we introduce three diagnostic-driven instantiations: the Human-inspired Supervised Adaptive Optimizer (HSAO), Hybrid Error-Diagnostic Reinforcement Learning (HED-RL) for actor-critic methods, and the Meta-Learned Learning Policy (MLLP). Under standard smoothness assumptions, we establish bounded effective updates and stability properties for all cases. Representative diagnostic illustrations in actor-critic learning highlight how the proposed signals modulate adaptation in response to TD error structure. Overall, this work elevates error evolution to a first-class object in adaptive learning and provides an interpretable, lightweight foundation for reliable learning in dynamic environments.
@@
adaptive learning; bias-noise-alignment; nonstationary environments; reinforcement learning; learned optimizers; stability; error evolution; interpretable diagnostics.
@@ 1 Introduction
Modern learning systems are increasingly deployed in environments that are nonstationary, noisy, and safety-critical, where the statistical properties of data and objectives evolve over time. In such settings, learning dynamics are often dominated not only by stochastic gradients but also by structured changes in operating conditions, feedback reliability, and temporal dependencies. Examples include sequential prediction under regime shifts, control under uncertain dynamics, and continual or task-adaptive learning. Despite remarkable progress in optimization and reinforcement learning (RL), ensuring stable, interpretable, and reliable adaptation under these conditions remains a fundamental challenge.
@@
Most widely used optimization algorithms adapt learning rates based on instantaneous or exponentially averaged gradient statistics [rubio2017convergence]. Methods such as Adam and AdamW [kingma2014adam, loshchilov2017decoupled, zaheer2018adaptive], large-batch optimizers such as LAMB [you2019large], sign-based methods such as Lion [chen2023lion], and second-order approximations such as Sophia [liu2023sophia] have proven effective across a wide range of applications. However, these methods primarily react to gradient magnitude and variance, without explicitly accounting for how the error signal itself evolves during training. In nonstationary settings, this limitation can lead to unstable convergence, oscillatory behavior, or overly conservative adaptation, particularly in long-horizon sequence models and control-oriented learning problems. Related challenges arise in RL, where training stability is strongly influenced by the structure of temporal-difference (TD) errors. Stabilization techniques such as trust-region methods [schulman2015trust], proximal policy optimization (PPO) [schulman2017proximal], entropy regularization, and advantage estimation reduce variance and catastrophic updates, but they do not explicitly distinguish between systematic error drift and stochastic fluctuations in the learning signal. As a result, policy learning remains highly sensitive to hyperparameters, noise, and implementation details, particularly under changing dynamics or partial observability [henderson2018deep].
@@
Meta-learning and learned optimizers provide another avenue for adaptation by learning update rules or initializations that generalize across tasks. Approaches such as Model-agnostic meta-learning (MAML) [finn2017model], Meta-SGD [li2017meta], Reptile, and recurrent learned optimizers [andrychowicz2016learning] demonstrate that optimization itself can be learned from data. While powerful, these methods typically condition on gradients and internal hidden states, offering limited interpretability of why or when adaptation occurs. Moreover, without explicit mechanisms to assess the reliability of the error signal, learned optimizers may still exhibit brittle behavior under unseen or highly nonstationary conditions. A common thread across supervised learning, RL, and meta-learning is that adaptation decisions are largely driven by gradient-based signals, while the temporal structure of the error signal is treated as incidental. In contrast, biological and human motor learning exhibit a qualitatively different behavior, adaptation is modulated based on perceived error trends, variability, and repeated overshoot rather than instantaneous feedback alone. Persistent drift triggers corrective action, high variability leads to cautious updates, and repeated overshoot prompts attenuation of movement amplitude. These observations suggest that error evolution contains interpretable structure that can serve as a principled control signal for adaptive learning.
@@
Motivated by this perspective, we propose a diagnostic-driven adaptive learning framework that explicitly models error dynamics through a simple yet expressive decomposition into three components: bias, capturing persistent error drift; noise, capturing stochastic variability around the trend; and alignment, capturing systematic overshoot arising from repeated update directions. These diagnostics are computed online from lightweight statistics of loss or TD-error evolution and are independent of model architecture or task domain. Rather than replacing gradient-based learning, the proposed diagnostics complement gradients by regulating update magnitude, direction, and exploration based on the reliability and structure of the learning signal. The central contribution of this work is to demonstrate that a bias-noise-alignment decomposition of error dynamics provides a unifying backbone for adaptive learning across multiple paradigms. We show that the same diagnostic signals can be systematically reused to stabilize supervised optimization, regulate actor-critic RL, and condition learned optimizers for rapid task adaptation. By explicitly elevating error evolution to a first-class object, this framework bridges ideas from optimization, control, and meta-learning within a single interpretable perspective. Under standard smoothness assumptions, we establish bounded effective updates and descent-style stability guarantees for diagnostic-driven adaptation. The proposed approach does not replace existing optimizers or control algorithms; rather, it enhances their robustness when learning dynamics become biased, noisy, or nonstationary. The major contributions of this work are summarized below.
@@
  • We introduce a lightweight bias-noise-alignment decomposition of error dynamics that is model-agnostic and computable online from loss or TD-error trajectories.
@@
  • We instantiate a single diagnostic framework across supervised learning, actor-critic RL, and meta-learned optimizers, demonstrating its role as a unifying mechanism for adaptive learning.
@@ 2 Review of Literature
This work connects adaptive optimization, RL stabilization, and meta-learning. We briefly review relevant literature in these areas and highlight the limitations that motivate a diagnostic-driven perspective on adaptive learning.
@@ 2.1 Adaptive Optimization Methods
Stochastic gradient-based optimization remains the backbone of modern machine learning. Classical methods such as SGD with momentum adapt update directions using first-order information, while adaptive methods such as Adam and AdamW [kingma2014adam, loshchilov2017decoupled] further normalize updates using gradient moments. These approaches provide robustness to scale differences across parameters and have become standard in training deep networks [ioannou2023adalip]. Extensions such as AdaFactor [shazeer2018adafactor], LAMB [you2019large], and Lion [chen2023lion] target improved scalability, memory efficiency, or stability in large-batch and large-model regimes. More recently, sharpness-aware optimization methods such as SAM [foret2020sharpness] and its variants have been proposed to improve generalization by explicitly seeking flat minima. While these methods introduce an implicit sensitivity to local loss geometry, they typically rely on inner maximization steps or gradient perturbations and do not explicitly model the temporal evolution of the error signal. Second-order or quasi-second-order approaches, including Sophia [liu2023sophia], approximate curvature information to improve convergence speed, but remain primarily reactive to instantaneous gradient statistics. A common limitation across these methods is that adaptation decisions are driven by gradient magnitude, variance, or curvature proxies, rather than by structured information about how the error evolves over time. As a result, these optimizers do not explicitly distinguish between persistent error drift, stochastic fluctuations, or repeated overshoot, which can lead to instability or overly conservative updates under nonstationary learning conditions.
@@ 2.2 Stability in Reinforcement Learning
RL introduces additional challenges due to delayed rewards, bootstrapping, and nonstationary data distributions induced by policy updates  [dulac2021challenges]. actor-critic methods mitigate variance by learning a value function, while advantage estimation techniques further stabilize gradient estimates. Trust-region approaches such as TRPO [schulman2015trust] constrain policy updates to avoid destructive changes, and PPO [schulman2017proximal] provides a computationally efficient approximation through clipped objectives. Entropy regularization is commonly used to encourage exploration and prevent premature convergence. However, entropy coefficients are typically fixed or scheduled heuristically, without explicit consideration of the structure of the TD error. Empirical studies have shown that deep RL methods remain highly sensitive to hyperparameters, random seeds, and implementation details [henderson2018deep], particularly under noisy or changing dynamics. While existing stabilization techniques limit update magnitude or variance, they do not explicitly distinguish whether instability arises from systematic bias in TD errors or from stochastic noise. Consequently, learning algorithms may respond uniformly to fundamentally different error regimes, motivating the need for diagnostic signals that can disentangle these effects.
@@ 2.3 Meta-Learning and Learned Optimizers
Meta-learning aims to enable rapid adaptation across tasks by learning how to learn. MAML [finn2017model] focuses on learning an initialization that supports fast task-specific adaptation using standard gradient descent. Extensions such as Meta-stochastic gradient descent (Meta-SGD) [li2017meta] jointly learn initialization and per-parameter learning rates, improving flexibility and adaptation speed. An alternative line of work treats the optimization process itself as a learning problem [chen2022learning]. Learned optimizers parameterize update rules using neural networks and are trained through meta-gradient descent across tasks [andrychowicz2016learning]. Recurrent and coordinate-wise learned optimizers have demonstrated impressive data efficiency in few-shot and continual learning settings. Despite these advances, learned optimizers often rely on opaque internal states and gradient inputs, making it difficult to interpret or reason about their adaptation behavior. Moreover, without explicit mechanisms to assess the reliability of the learning signal, learned optimizers may generalize poorly to tasks with substantially different noise characteristics or nonstationary dynamics. This limitation suggests that conditioning learned updates on interpretable diagnostics of error structure may improve robustness and transparency.
@@ 2.4 Human-Inspired and Control-Theoretic Perspectives
A broad class of learning algorithms has been motivated by principles drawn from human cognition, motor learning, and classical control theory. Representative examples include gain scheduling and adaptive step-size control in optimization, uncertainty-aware exploration and entropy modulation in RL, and curriculum or experience-driven adaptation in meta-learning. These approaches have demonstrated empirical benefits in stabilizing training and improving convergence in complex, nonstationary environments. Despite their success, most existing human-inspired methods incorporate adaptivity implicitly or heuristically. In practice, adaptation is often driven by gradient magnitude, loss curvature, or hand-designed schedules, without an explicit characterization of the reason behind the unstable learning. From a control-theoretic perspective, however, instability is rarely monolithic. Persistent bias such as systematic drift, stochastic noise such as measurement or process variability, and repeated directional excitation such as overshoot or resonance are treated as distinct phenomena, each requiring different compensatory mechanisms. When these effects are conflated, corrective actions may be either overly aggressive or excessively conservative. A similar limitation is observed across adaptive optimization, RL stabilization, and meta-learning. Although these methods provide powerful tools for learning in high-dimensional and uncertain settings, they remain largely gradient-centric and do not explicitly model the temporal structure of error evolution. In particular, they do not distinguish between systematic and stochastic sources of instability, nor do they provide an interpretable signal for modulating update magnitude, direction, or exploration based on error reliability.
The present work addresses this gap by introducing a diagnostic decomposition of error dynamics into bias, noise, and alignment components. Inspired by both human motor learning and control-theoretic reasoning, these diagnostics explicitly separate persistent drift, stochastic variability, and directional overshoot phenomena that humans naturally compensate for by adjusting step size, movement direction, and caution under uncertainty. Rather than serving as ad hoc heuristics, the proposed diagnostics form a unified and interpretable control signal that regulates learning across supervised optimization, actor-critic RL, and meta-learned optimizers. In this sense, the framework operationalizes human-inspired adaptation through a principled, task-agnostic representation of error evolution.
@@ 3 Diagnostic Decomposition of Error Dynamics
This section discusses the diagnostic framework used in this study to explicitly model the temporal evolution of error during learning [lewkowycz2020training]. Rather than relying solely on instantaneous gradients, we extract interpretable signals from the dynamics of the loss or TD error and use them as lightweight controls for adaptive learning. The proposed decomposition separates error dynamics into three complementary components namely, bias, noise, and alignment. These diagnostics are general, model-agnostic, and applicable across supervised learning, RL, and meta-learning. The details of each mechanism are discussed in the section below.
@@ 3.1 Error Evolution Signal
Let $\ell_{t}$ denote a scalar learning signal evaluated at iteration $t$. In supervised learning, $\ell_{t}$ may correspond to a training or validation loss, while in RL it may represent a TD error or a batch-level surrogate thereof. Rather than treating $\ell_{t}$ as an isolated quantity, we focus on its temporal evolution. We define the incremental error signal which captures whether learning progress is improving, stagnating, or degrading over time. Unlike raw gradients, $\Delta\ell_{t}$ directly reflects changes in performance and aggregates the effects of optimization, stochasticity, and model mismatch.